OpenStudy (calculusxy):

In slope-intercept form what is the line that is perpendicular to 4x + 2y = 9 passing thru (-3, 4)

OpenStudy (calculusxy):

@Nnesha

OpenStudy (calculusxy):

I got \[\large y = -\frac{1}{2}x + \frac{9}{2}\]

Nnesha (nnesha):

hmmm what was your first step ?? how did you get 1 at the numerator

OpenStudy (calculusxy):

I found out the slope

Nnesha (nnesha):

ohh i see do you mean that's the equation of perpendicular line ??

OpenStudy (calculusxy):

I am not sure that's what the question says

Nnesha (nnesha):

oh i thought u rewrite the first equation in slope intercept nvm wait a sec let me check it

Nnesha (nnesha):

perpendicular lines would have the `different ` y-intercept slope is correct how did you get 9/2 for y-intercept ?

OpenStudy (calculusxy):

`2y = 9 - 4x` \(\leftarrow\) that's what I got after moving the 4x to the right side of equal sign. To isolate \(y\) I divide both sides by \(2\) and then got \(\frac{9}{2}\).

Nnesha (nnesha):

alright 9/2 is y-intercept of 4x+2y=9 equation for perpendicular equation we need to sub given point which is (-3,4)

Nnesha (nnesha):

\[\rm y=-\frac{1}{2}x+b\] this is our new equation substitute x and y for (-3,4) to find y-intercept

OpenStudy (tommynaut):

There are a few steps to this problem. 1. Rearrange your original equation from the form y = mx + b. The whole point of that was to find the gradient of your original equation, m. 2. Now, you know the gradient of your perpendicular equation will be -1/m, right? 3. Use your new perpendicular gradient, along with the point given in the question, and plug them into the point-slope formula.

OpenStudy (tommynaut):

into the form y = mx + b, I mean

OpenStudy (calculusxy):

\[4= -\frac{ 1 }{ 2 }(-3) + b \rightarrow 4 = \frac{3}{2} + b \rightarrow b = \frac{5}{2}\]

Nnesha (nnesha):

looks good that's the y-intercept of perpendicular line

OpenStudy (calculusxy):

so our slope-int form would be: \[\large \tt y = -\frac{1}{2}x + \frac{5}{2}\]

Nnesha (nnesha):

yes right parallel lines = same slope different y-intercept perpendicular lines = slope(negative reciprocal )

OpenStudy (calculusxy):

Thanks!

OpenStudy (calculusxy):

wait.. wouldn't the slope be 1/2x instead of -1/2x?

OpenStudy (calculusxy):

@Nnesha

Nnesha (nnesha):

opps ye that's correct first one is negative so 2nd one would be positive sorry i should mentioned that

OpenStudy (calculusxy):

no prob :)